# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 2. Collision Trees
The main ingredient to prove Theorem 1.1 is the representation of the dynamics through collision trees. This will allow us to estimate the difference between the idealised behaviour as predicted by the Boltzmann equation and the empirical, genuine particle dynamics. We use the same structures as in [MT10, MT12] developed for nonlinear gainless dynamics. In [MST18, MS18] these were adapted to a Rayleigh gas of hard-sphere dynamics, such that collision trees have height 2 and essentially become lists. Very detailed proofs and explanations in the Rayleigh cases are given in [Sto17].
@@
where $v_{j}$, $j=1,...,n(\Phi)$ is the velocity of each background particle in the collision history $\Phi$ and $v(s),$ $s\in[0,T]$ is the velocity of the tagged particle in the same history for different times.
@@
$\begin{cases}\partial_{t}P_{t}(\Phi)&=c[Q_{t}^{+}[P_{t}](\Phi)-P_{t}(\Phi)Q^{-}_{\tau}(\Phi)],\\ P_{0}(\Phi)&=f_{0}(x_{0},v_{0})\mathds{1}_{n(\Phi)=0},\end{cases}$ (2.1)
@@
$Q_{t}^{+}[P_{t}](\Phi)=\begin{cases}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)&n(\Phi)>0,\\ 0&n(\Phi)=0,\end{cases}$ (2.2)
@@
$Q^{-}_{\tau}(\Phi)=\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}{\omega}\,\mathrm{d}\bar{v},$ (2.3)
@@
A collision history $\Phi\in\mathcal{MT}$ is called re-collision free at diameter $\varepsilon$ if for all $j=1,...,n(\Phi)$ and for all $t\in[0,T]\setminus(t_{j},t_{j}+\tau*)$,
$|x(t)-x_{j}(t)|>\varepsilon.$
That is, if the tagged particle and a background particle $j$ collide at time $(t_{j},t_{j}+\tau^{*})$ then the tagged particle has not previously collided and will not re-collide with the background particle $j$ up to time $T$. Define the set
@@
Here $\tau^{*}$ is the scattering time, i.e., the time it takes for the collision to take place, in the short range potential model.
@@
$\mathcal{G}(\varepsilon):=\{\Phi\in\mathcal{MT}:\ n(\Phi)\leq M(\varepsilon),\ \mathcal{V}(\Phi)<V(\varepsilon),\ \Phi\in R(\varepsilon)\cap S(\varepsilon)\ \mathrm{and}\ \Phi\text{ is\ non-}\mathrm{grazing}\},$
for any decreasing functions $V,M:(0,\infty)\to[0,\infty)$ such that $\lim_{\varepsilon\to 0}V(\varepsilon)=\lim_{\varepsilon\to 0}M(\varepsilon)=\infty.$
3. The collision kernel $B$
3.1. Deriving the collision kernel $B$
We will require a detailed understanding of the collision kernel $B$ in (1.4). In [GSRT13, Section 8.3] an implicit formula for the collision kernel $B$ is derived, the process is as below. First, for fixed $x_{1}$ they calculate the surface measure on the sphere $\mathbb{S}^{d-1}:=\{y\in\mathds{R}^{d}\ :\ |y-x_{1}|=\varepsilon\}$ which is defined by $d\sigma_{1}$ and also $x_{2}\in\mathbb{S}^{d-1}$. Here $\varepsilon$ is the radius of the sphere. Now parameterise the sphere by ($\alpha,\psi$) $\in[0,\pi]\times[0,2\pi]$. Also, $\alpha$ is the angle between $\delta v=v_{1}-v_{2}$ and $\delta x=x_{1}-x_{2}$. Then
$\mathrm{d}\sigma_{1}=\varepsilon^{d-1}(\sin\alpha)^{d-2}\mathrm{d}\alpha\mathrm{d}\psi.$
@@ 3.1. Deriving the collision kernel $B$
$\mathrm{d}\omega=(\sin\Theta)^{d-2}\mathrm{d}\Theta\mathrm{d}\psi.$ (3.1)
@@
$\frac{1}{\varepsilon}(x_{1}-x_{2})\cdot(v_{1}-v_{2})\mathrm{d}\sigma_{1}=\frac{1}{\varepsilon}\varepsilon|v_{1}-v_{2}|\cos\alpha\varepsilon^{d-1}(\sin\alpha)^{d-2}\mathrm{d}\alpha\mathrm{d}\psi$
$=$ $\varepsilon^{d-1}|v_{1}-v_{2}|\cos\alpha(\sin\alpha)^{d-2}\mathrm{d}\alpha\mathrm{d}\psi=\varepsilon^{d-1}|v_{1}-v_{2}|\mathcal{J}_{0}^{d-2}\mathrm{d}\mathcal{J}_{0}\mathrm{d}\psi$
$=$ $\varepsilon^{d-1}|v_{1}-v_{2}|\mathcal{J}_{0}^{d-2}\partial_{\Theta}\mathcal{J}_{0}\mathrm{d}\Theta\mathrm{d}\psi.$
@@
$\mathrm{d}\Theta\,\mathrm{d}\psi=(\sin\Theta)^{2-d}\mathrm{d}\omega$
@@
$\frac{1}{\varepsilon}(x_{1}-x_{2})\cdot(v_{1}-v_{2})\mathrm{d}\sigma_{1}=\varepsilon^{d-1}|v_{1}-v_{2}|\mathcal{J}_{0}^{d-2}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{2-d}\mathrm{d}\omega.$
These calculations are local in the scattering angle $\Theta$, which allows us include the angular cut-off.
@@
The scattering cross-section with cut-off $\sigma$ is defined for $|v_{1}-v_{2}|>0$ and $\Theta\in(0,\frac{\pi}{2}-\sigma)$, by $\mathcal{J}_{0}^{d-2}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{2-d}$. Therefore, the collision kernel $B$ is defined by
$B(v_{1}-v_{2},\Theta):=\begin{cases}|v_{1}-v_{2}|\mathcal{J}_{0}^{d-2}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{2-d}&\mbox{ if }\Theta\in(0,\frac{\pi}{2}-\sigma),\\ 0&\mbox{ if }\Theta\in[\frac{\pi}{2}-\sigma,\frac{\pi}{2}].\end{cases}$ (3.2)
By abuse of notation we may write $B(v_{1}-v_{2},\Theta)=B(v_{1}-v_{2},\omega)$. Here, $|v_{1}-v_{2}|$ is the kinetic part of the collision kernel and the remaining is the angular part of the collision kernel. The collision kernel $B$ for $d=3$ becomes
$B(v_{1}-v_{2},\Theta)=\begin{cases}|v_{1}-v_{2}|\mathcal{J}_{0}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{-1}&\mbox{ if }\Theta\in(0,\frac{\pi}{2}-\sigma),\\ 0&\mbox{ if }\Theta\in[\frac{\pi}{2}-\sigma,\frac{\pi}{2}].\end{cases}$ (3.3)
3.2. Formula of the collision kernel $B$
This subsection is devoted in finding the formula of the collision kernel $B$ for two particles that are interacting via the short range potential introduced in the particle dynamics subsection in the introduction. We are doing so by finding the inverse of the function $\Theta$ with respect to $\mathcal{J}_{0}$ away from the value $\Theta=\frac{\pi}{2}$. That is because, near this value of $\Theta$, the collision kernel is not integrable due to grazing collisions. Then we apply the above Definition 3.1 for these values of $\Theta$ and for large relative velocities. The main proposition of this section is the following.
@@ 3.2. Formula of the collision kernel $B$
For any $n\in(2,5]$, the collision kernel $B$ is bounded for large relative velocities $|v_{1}-v_{2}|>0$ and for the deviation angle $\Theta\in(0,\frac{\pi}{2}-\sigma)$, $\sigma>0$, $B$ takes the form
$B(v_{1}-v_{2},\Theta)$ $=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\bigg{[}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\bigg{]}\Theta(\sin\Theta)^{-1},$ (3.4)
where $a$ is a fixed number and $K>0$. Furthermore, it holds true that
$\bigg{|}B(v_{1}-v_{2},\Theta)-\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\bigg{|}$ $=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})$
@@
In order to give a proof of this proposition, we first need to prove Proposition 3.4 which is about finding a formula of the function $\mathcal{J}_{0}$ with respect to $\Theta$. Then we can compute the derivative of $\mathcal{J}_{0}$ with respect to $\Theta$ and then use (3.3) to give the formula (3.4) of the collision kernel $B$.
@@
(Upper and Lower bound for the collision kernel $B$) For $n\in(2,5]$, large $|v_{1}-v_{2}|>0$ and $\Theta\in(0,\frac{\pi}{2}-\sigma)$, we have the following bounds
$\frac{1}{2}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\Theta(\sin\Theta)^{-1}$ $\leq B(v_{1}-v_{2},\Theta)\leq 3\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\Theta(\sin\Theta)^{-1}.$ (3.5)
@@
Immediate by the way that the collision kernel $B$ is defined. ∎
@@
For any $n\in(2,5]$, for large relative velocities $|v_{1}-v_{2}|>0$ and for $\Theta\in(0,\frac{\pi}{2}-\sigma)$, $\sigma>0$ the formula of $\mathcal{J}_{0}(\Theta)$ is given by
$\mathcal{J}_{0}(\Theta)=\frac{(2K)^{\frac{1}{n-1}}}{a}\bigg{(}1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\bigg{)}\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}},$ (3.6)
where $\mathcal{E}_{0}=|v_{1}-v_{2}|^{2}$, $a$ is a fixed real number and $K>0$.
@@
$\Theta=\arcsin{J_{0}}+\int_{\lambda}^{x_{0}}\frac{\mathrm{d}x}{\sqrt{1-x^{2}-(\frac{x}{b})^{n-1}}}$ (3.7)
@@
$1-x_{0}^{2}-\bigg{(}\frac{x_{0}}{b}\bigg{)}^{n-1}=0,\quad\lambda=J_{0}\bigg{(}1+\frac{2K}{\mathcal{E}_{0}}\bigg{)}^{\frac{1}{2}},\quad b=J_{0}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}},$
@@
$\mathcal{J}_{0}:=\frac{|(x_{1}-x_{2})\times(v_{1}-v_{2})|}{\varepsilon|v_{1}-v_{2}|}:=\sin\alpha,$
@@
$\Theta=\arcsin{J_{0}}+\int_{\frac{\lambda}{b}}^{u_{0}}\frac{b\,\mathrm{d}u}{\sqrt{1-b^{2}u^{2}-u^{n-1}}}$
@@
$\Theta=\arcsin{J_{0}}+\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{bu_{0}\,\mathrm{d}y}{\sqrt{1-b^{2}u_{0}^{2}y^{2}-(u_{0}y)^{n-1}}}$
@@
$\Theta=\arcsin{J_{0}}+\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{bu_{0}\,\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}}.$
@@
$J_{0}^{2}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{2}{n-1}}u_{0}^{2}=1-u_{0}^{n-1},$ (3.9)
@@
$\Sigma(\mathcal{J}_{0}):=\arcsin{J_{0}}+\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{bu_{0}\,\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}}$
@@
$\Sigma^{\prime}(\mathcal{J}_{0})=\frac{1}{\sqrt{1-J_{0}^{2}}}$ $+\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}u_{0}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}}$
$+J_{0}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}u_{0}\frac{\mathrm{d}}{\mathrm{d}\mathcal{J}_{0}}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}},$
@@
$\frac{\mathrm{d}}{\mathrm{d}\mathcal{J}_{0}}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}}$ $=-\frac{-\frac{\lambda}{b}\frac{1}{u_{0}^{2}}\overbrace{\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}}^{=0}}{\sqrt{1-(1-u_{0}^{n-1})(\frac{\lambda}{bu_{0}})^{2}-(\frac{\lambda}{b})^{n-1}}}$
$\quad-\frac{1}{2}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{(n-1)u_{0}^{n-2}(y^{2}-y^{n-1})\overbrace{\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}}^{=0}\mathrm{d}y}{(1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1})^{\frac{3}{2}}}$
$=0.$
Here we used that the derivative of $u_{0}$ with respect to $\mathcal{J}_{0}$ is zero, see Lemma 3.5 below. Therefore the derivative of $\Sigma(\mathcal{J}_{0})$ with respect to $\mathcal{J}_{0}$ is
$\Sigma^{\prime}(\mathcal{J}_{0})=\frac{1}{\sqrt{1-J_{0}^{2}}}$ $+\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}u_{0}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}}.$ (3.10)
The second term, for $u_{0}=1$ becomes
$\underbrace{\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}}_{=\frac{b}{\mathcal{J}_{0}}}\int_{\frac{\lambda}{b}}^{1}\frac{\mathrm{d}y}{\sqrt{1-y^{n-1}}}.$ (3.11)
By using the change of variable $w:=1-y$ the integral becomes

### v2

@@ 2. Collision Trees
The main ingredient to prove Theorem 1.1 is the representation of the dynamics through collision trees. This will allow us to estimate the difference between the idealised behaviour as predicted by the Boltzmann equation and the empirical, genuine particle dynamics. We use similar structures as in [MT10, MT12] developed for nonlinear gainless dynamics. In [MST18, MS18] these were adapted to a Rayleigh gas of hard-sphere dynamics, such that collision trees have height 2 and essentially become lists. Very detailed proofs and explanations in the Rayleigh cases are given in [Sto17].
@@
where $v_{j}$, $j=1,...,n(\Phi)$ is the velocity of each background particle in the collision history $\Phi$ and $v(s),$ $s\in[0,T]$ is the velocity of the tagged particle in the same history for different times. We denote the smallest relative speed of the $n(\Phi)$ collisions as
$\iota(\Phi):=\min_{j=1,...,n(\Phi)}|v(t_{j}^{-})-v_{j}|.$
@@
$\begin{cases}\partial_{t}P_{t}(\Phi)&=c[Q_{t}^{+}[P_{t}](\Phi)-Q^{-}_{t}[P_{t}](\Phi)],\\ P_{0}(\Phi)&=f_{0}(x_{0},v_{0})\mathds{1}_{n(\Phi)=0},\end{cases}$ (2.1)
@@
$Q_{t}^{+}[P_{t}](\Phi)=\begin{cases}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)&n(\Phi)>0,\\ 0&n(\Phi)=0,\end{cases}$ (2.2)
@@
$Q^{-}_{t}[P_{t}](\Phi)=P_{t}(\Phi)\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau)-\bar{v},\omega)\,\mathrm{d}{\omega}\,\mathrm{d}\bar{v},$ (2.3)
@@
A collision history $\Phi\in\mathcal{MT}$ is called re-collision free at diameter $\varepsilon$ if for all $j=1,...,n(\Phi)$ and for all $t\in[0,T]\setminus(t_{j},t_{j}+\tau_{j}^{*})$,
$|x_{0}(t)-x_{j}(t)|>\varepsilon.$
That is, if the tagged particle and a background particle $j$ collide at time $(t_{j},t_{j}+\tau_{j}^{*})$ then the tagged particle has not previously collided and will not re-collide with the background particle $j$ up to time $T$. Define the set
@@
Here $\tau_{j}^{*}$ is the scattering time, i.e., the time it takes for the collision to take place, in the short range potential model. We also define the maximal collision time in a tree as
$(\Phi)=\max_{j=1,...,n(\Phi)}\tau_{j}^{*}.$
Definition 2.6.
A collision history $\Phi\in\mathcal{MT}$ is called binary at diameter $\varepsilon$ if for all $j=1,...,n(\Phi)$ the collision times $(t_{j},t_{j}+\tau_{j}^{*})$ are pairwise disjoint.
That is, the tagged particle will only interact with at most a single background particle at any given time. Define the set
$B(\varepsilon):=\{\Phi\in\mathcal{MT}:\ \Phi\text{\ is\ binary\ at\ diameter}\ \varepsilon\}.$
@@
$\mathcal{G}(\varepsilon):=\{\Phi\in\mathcal{MT}:\ n(\Phi)\leq M(\varepsilon),\ \mathcal{V}(\Phi)<$ $V(\varepsilon),\ \vartheta(\Phi)>I(\varepsilon),\ \iota(\Phi)>J(\Phi),$
[…]
for any strictly increasing $I,J:(0,\infty)\to[0,\infty)$ such that $\lim_{\varepsilon\to 0}I(\varepsilon)=\lim_{\varepsilon\to 0}J(\varepsilon)=0$ and any decreasing functions $V,M:(0,\infty)\to[0,\infty)$ such that $\lim_{\varepsilon\to 0}V(\varepsilon)=\lim_{\varepsilon\to 0}M(\varepsilon)=\infty.$
The functions $I$, $J$, $V$ and $M$ will be defined in subsection 4.9.
3. The collision kernel $\mathcal{B}$
3.1. Deriving the collision kernel $\mathcal{B}$
We will require a detailed understanding of the collision kernel $\mathcal{B}$ in (1.3). In [GSRT13, Section 8.3] an implicit formula for the collision kernel $\mathcal{B}$ is derived, the process is as below. First, for fixed $x_{1}$ they calculate the surface measure on the sphere $\{y\in\mathds{R}^{3}\ :\ |y-x_{1}|=\varepsilon\}$ which is defined by $d\sigma_{1}$ and also $x_{2}$ belongs in this sphere. Here $\varepsilon$ is the radius of the sphere. Now parametrise the sphere by ($\alpha,\psi$) $\in[0,\pi]\times[0,2\pi]$. Also, $\alpha$ is the angle between $\delta v=v_{1}-v_{2}$ and $\delta x=x_{1}-x_{2}$. Then
$\mathrm{d}\sigma_{1}=\varepsilon^{2}\sin\alpha\mathrm{d}\alpha\mathrm{d}\psi.$
@@ 3.1. Deriving the collision kernel $\mathcal{B}$
$\mathrm{d}\omega=\sin\Theta\mathrm{d}\Theta\mathrm{d}\psi.$ (3.1)
@@
[…]
@@
$\mathrm{d}\Theta\,\mathrm{d}\psi=(\sin\Theta)^{-1}\mathrm{d}\omega$
@@
$\frac{1}{\varepsilon}(x_{1}-x_{2})\cdot(v_{1}-v_{2})\mathrm{d}\sigma_{1}=\varepsilon^{2}|v_{1}-v_{2}|\mathcal{J}_{0}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{-1}\mathrm{d}\omega.$
@@
The scattering cross-section is defined for $|v_{1}-v_{2}|>0$ by $\mathcal{J}_{0}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{-1}$. Therefore, the collision kernel $\mathcal{B}$ is defined by
$\mathcal{B}(v_{1}-v_{2},\Theta):=$ $|v_{1}-v_{2}|\mathcal{J}_{0}\partial_{\Theta}\mathcal{J}_{0}(\sin\Theta)^{-1}.$ (3.2)
By abuse of notation we may write $\mathcal{B}(v_{1}-v_{2},\Theta)=\mathcal{B}(v_{1}-v_{2},\omega)$. Here, $|v_{1}-v_{2}|$ is the kinetic part of the collision kernel and the remaining is the angular part of the collision kernel.
3.2. Formula of the collision kernel $\mathcal{B}$
This subsection contains computations of finding the formula of the collision kernel $\mathcal{B}$ for two particles that are interacting via the short range potential introduced in the particle dynamics subsection in the introduction. We are doing so by finding the inverse of the function $\Theta$ with respect to $\mathcal{J}_{0}$. Then we apply the above Definition 3.1 and analyse the behaviour for large relative velocities. We start by observing that $\mathcal{B}$ bounded for bounded velocities
Lemma 3.2.
For any $n\in(3,5]$, the collision kernel $\mathcal{B}$ is bounded on bounded sets for $(v_{1}-v_{2},\Theta)\in\mathds{R}^{3}\times[0,\pi/2]$.
Proof.
Following [GSRT13][Lemma 8.3.1], we observe that for fixed $|v_{1}-v_{2}|$, that $\mathcal{J}_{0}(\Theta)$ and $\partial_{\Theta}\mathcal{J}_{0}(\Theta)$ are well-defined, bounded and continuous for all $\Theta\in[0,\pi/2]$, such that $\mathcal{J}_{0}(0)=0$ and $\partial_{\Theta}\mathcal{J}_{0}(0)<\infty$, which implies continuity of $\mathcal{B}$ for $|v_{1}-v_{2}|>0$. A simple inspection of the for $\Theta(|v_{1}-v_{2}|^{2},\mathcal{J}_{0})$ and $\frac{\partial\Theta}{\partial\mathcal{J}_{0}}(|v_{1}-v_{2}|^{2},\mathcal{J}_{0})$ yields that in the limit $|v_{1}-v_{2}|^{2}\to 0$, the terms simplify to $\mathcal{J}_{0}(\Theta)=\sin(\Theta)$ and $\partial_{\Theta}\mathcal{J}_{0}(\Theta)=\cos(\Theta).$ Hence $\mathcal{B}$ as given in (3.2) for our interaction potentials is also continuous at $|v_{1}-v_{2}|=0$ such that it is bounded for bounded relative velocities $|v_{1}-v_{2}|$ uniformly in $\Theta\in[0,\pi/2]$. ∎
The main proposition of this section is the following.
@@ 3.2. Formula of the collision kernel $\mathcal{B}$
For any $n\in(3,5]$, the collision kernel $\mathcal{B}$ is bounded for large relative velocities $|v_{1}-v_{2}|\to\infty$ and $\mathcal{B}$ takes the form
$\mathcal{B}(v_{1}-v_{2},\Theta)$ $=\frac{2^{\frac{2}{n-1}}}{a^{2}}\bigg[|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\bigg]\Theta(\sin\Theta)^{-1},$ (3.3)
where $a$ is a fixed number. Furthermore, it holds true that
$\bigg|\mathcal{B}(v_{1}-v_{2},\Theta)-\frac{2^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\bigg|$ $=\frac{2^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})$
@@
We note here that when we use the Landau notation in (3.3) for large relative velocities, we mean $\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})$, as $|v_{1}-v_{2}|\to\infty$ where we mean the standard Landau notation $f=\mathcal{O}(g)$ as $x\to\infty$ if there exists a constant $C$ such that $|f(x)|\leq C|g(x)|$ for all $x$ sufficiently large. The Proposition immediately implies upper and lower bounds.
@@
(Upper and Lower bound for the collision kernel $\mathcal{B}$) For $n\in(3,5]$ and large $|v_{1}-v_{2}|>0$ we have the following bounds.
$\frac{1}{2}\frac{2^{\frac{2}{n-1}}}{a^{2}}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\Theta(\sin\Theta)^{-1}$ $\leq\mathcal{B}(v_{1}-v_{2},\Theta)\leq 3\frac{2^{\frac{2}{n-1}}}{a^{2}}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\Theta(\sin\Theta)^{-1}.$ (3.4)
@@
Immediate by the way that the collision kernel $\mathcal{B}$ is defined. ∎
In order to give a proof of this proposition, we first need to prove Proposition which is about finding a formula of the function $\mathcal{J}_{0}$ with respect to $\Theta$. Then we can compute the derivative of $\mathcal{J}_{0}$ with respect to $\Theta$ and then we use (3.2) to give the formula (3.3) of the collision kernel $\mathcal{B}$. We first give the proof of a lemma needed for Proposition 3.6.
Lemma 3.5.
Proof.
With (3.8) we define
$F(u_{0},\mathcal{J}_{0}):=J_{0}^{2}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{2}{n-1}}u_{0}^{2}+u_{0}^{n-1}-1.$
We observe that $F(1,0)=0$ and
$\frac{\partial F}{\partial u_{0}}(u_{0},\mathcal{J}_{0})=2J_{0}^{2}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{2}{n-1}}u_{0}+(n-1)u_{0}^{n-2},$
so $\frac{\partial F}{\partial u_{0}}(1,0)=n-1\neq 0$, for $n\neq 1.$ Also
$\frac{\partial F}{\partial\mathcal{J}_{0}}(u_{0},\mathcal{J}_{0})=2J_{0}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{2}{n-1}}u_{0}^{2},$
which is equivalent to
$\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}\bigg|_{\mathcal{J}_{0}=0}=-\bigg(\frac{\partial F}{\partial u_{0}}(1,0)\bigg)^{-1}\frac{\partial F}{\partial\mathcal{J}_{0}}(1,0)=0,$
as required. ∎
@@
For any $n\in(3,5]$ and for large relative velocities $|v_{1}-v_{2}|>0$, the formula of $\mathcal{J}_{0}(\Theta)$ is given by
$\mathcal{J}_{0}(\Theta)=\frac{2^{\frac{1}{n-1}}}{a}\bigg(1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\bigg)\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}},$ (3.5)
where $\mathcal{E}_{0}=|v_{1}-v_{2}|^{2}$ and $a$ is a fixed real number.
@@
[…]
@@
$1-x_{0}^{2}-\bigg(\frac{x_{0}}{b}\bigg)^{n-1}=0,\quad\lambda=J_{0}\bigg(1+\frac{2}{\mathcal{E}_{0}}\bigg)^{\frac{1}{2}},\quad b=J_{0}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}},$
@@
$\mathcal{J}_{0}:=\frac{|(x_{1}-x_{2})\times(v_{1}-v_{2})|}{\varepsilon|v_{1}-v_{2}|}=:\sin\alpha,$
@@
[…]
@@
[…]
@@
[…]
@@
$J_{0}^{2}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{2}{n-1}}u_{0}^{2}=1-u_{0}^{n-1},$ (3.8)
@@
[…]
@@
[…]
$+J_{0}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}}u_{0}\frac{\mathrm{d}}{\mathrm{d}\mathcal{J}_{0}}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}},$ (3.9)
@@
$\frac{\mathrm{d}}{\mathrm{d}\mathcal{J}_{0}}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{\mathrm{d}y}{\sqrt{1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1}}}$ $=-\frac{-\frac{\lambda}{b}\frac{1}{u_{0}^{2}}\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}}{\sqrt{1-(1-u_{0}^{n-1})(\frac{\lambda}{bu_{0}})^{2}-(\frac{\lambda}{b})^{n-1}}}$
$\quad-\frac{1}{2}\int_{\frac{\lambda}{bu_{0}}}^{1}\frac{(n-1)u_{0}^{n-2}(y^{2}-y^{n-1})\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}\mathrm{d}y}{(1-(1-u_{0}^{n-1})y^{2}-(u_{0}y)^{n-1})^{\frac{3}{2}}}.$
Therefore, the derivative of $\Sigma(\mathcal{J}_{0})$ for $\mathcal{J}_{0}=0$ is
[…]
where we used that when $\mathcal{J}_{0}=0$, then $u_{0}=1$, by (3.8) and the fact that the derivative of $u_{0}$ with respect to $\mathcal{J}_{0}$ is zero, for $\mathcal{J}_{0}=0$, see Lemma 3.5 below. By using the change of variable w:=1yw:=1-y the above integral becomes